# How connection probability shapes fluctuations of neural population dynamics

### v1

@@ III.2 Second-order mesoscopic mean-field dynamics
Here, we have used that, for $\sigma=0$, the population transfer function $F(\bar{h},\sigma^{2})$ simplifies to the transfer function of a single Poisson neuron $\phi(\bar{h})=F(\bar{h}(t),0)$ and that $G(\bar{h},0)=0$. The 1st-order MF model (21) precisely corresponds to the limit $p\to 1$ while keeping the network size $N=C/p$ and the coupling strength $w=JpN$ constant. In the following, the 1st-order MF theory can thus always be obtained as a limit case of the 2nd-order theory by first re-expressing any occurrence of the parameters $C$ and $J$ by $pN$ and $w/(pN)$, respectively, and then taking the limit $p\to 1$.
Apart from a Gaussian approximation for large $N$, the 1st-order MF model is essentially equivalent to the mean-connectivity network, Eq. (5). Indeed, the 1st-order MF model is recovered from the exact mesoscopic mean-field dynamics, Eq. (7), of the mean-connectivity network, if the Gaussian approximation Eq. (14) is applied to the Poisson noise in Eq. (7). By comparing simulations of the microscopic models (quenched and annealed networks) with the 1st-order and 2nd-order MF models, we will thus be able to judge the mean-connectivity approximation used in previous studies [49, 35, 50] and the correction for the effect of random dilution ($p<1$) given by the 2nd-order MF theory.
In the limit $N\to\infty$, the 2nd-order mean-field model, Eq. (20), converges to either a one- or a two-dimensional, deterministic model depending on whether the connectivity is dense or sparse. In the dense limit, where $p$ is kept constant and, hence, $C=pN\to\infty$, the variance $\sigma^{2}$ as well as the fluctuations $N^{-1/2}\eta$ and $N^{-1/2}\xi$ vanish leaving a one-dimensional, deterministic firing rate model. This case is equivalent to the large-$N$ limit of the 1st-order MF model. In contrast, in the sparse limit, where $C$ is kept constant and, hence, $p=C/N\to 0$, the noise terms $N^{-1/2}\eta$ and $N^{-1/2}\xi$ vanish but the variance $\sigma^{2}$ does not. The result is a two-dimensional, deterministic mean-field model for the mean and variance of the input potentials similar to the classical model of Amari [22].
@@ III.2.1 Functions $F$ and $G$ for an error-function nonlinearity
being the effective steepness of the sigmoidal nonlinearity at the population level. Because $\beta_{\text{eff}}(\sigma^{2})\leq\beta_{\text{eff}}(0)=\beta$, the effective nonlinearity that governs the population dynamics has a reduced steepness compared to the the single-neuron nonlinearity (Fig. 3) and this reduction is caused by the strictly positive spread $\sigma>0$ of the membrane potentials in a non-fully connected random network.
@@ IV Analysis of the mesoscopic model
The reduction of the microscopic network model to a mesoscopic MF model in form of the system Eq. (20) allows us to calculate and analyze the first- and second-order statistics of mesoscopic quantities such as the population activity $A_{N}(t,\Delta t)$, the stochastic population rate $r(t)$ and the population-averaged input potential $\bar{h}(t)$.
@@ IV.1 Stationary mean population activity
Figure 3: Mean stationary population rate of the quenched and annealed networks and fixed-point solutions of the noiseless mean-field dynamics. (a) Graphical solution of the fixed point equation (25). Solid lines: function $F(\bar{h},\sigma^{2}(\bar{h}))$ for the 1st-order (orange, $\sigma^{2}(\bar{h})\equiv 0$) and 2nd-order mean-field model (blue, $\sigma^{2}(\bar{h})$ as in Eq. (25)). Black dashed line: left-hand side of Eq. (25). Cyan dash-dotted line: $F$ with the second argument fixed at the fixed-point value $\sigma_{0}^{2}$. Mean population rates as a function of (b) the external input $\mu_{0}$, (c) the connection probability $p$, and (d) the coupling strength $w$. Orange (blue) line: 1st-order (2nd-order) MF theory; red squares (green circles): annealed (quenched) network simulation; dotted line: analytical approximation (26). The inset in (d) shows the same data, but for a larger range of $w$ in a double logarithmic scaling. Grey dashed line: 2nd-order theory $r_{\infty}$ for the firing rate in the limit $w\to-\infty$ (Eq. (29)). Parameters: $N=1000$, $\beta=5$ mV-1 (a): $p=0.1$, $w=-0.7$ mVs, and $\mu_{0}=5$ mV. (b): $p=0.1$ and $w=-1$ mVs, (c): $\mu_{0}=10$ mV, $w=-1$ mVs, (d): $p=0.1$, $\mu_{0}=10$ mVs.
@@
We first analyze the fixed-point solutions $\bar{h}(t)=\bar{h}_{0}$, $\sigma^{2}(t)=\sigma_{0}^{2}$ and $\xi(t)=0$ of the deterministic system when the noise terms in Eq. (20) are set to zero and the external drive is constant, $\mu(t)=\mu_{0}$. In the following, we assume an inhibitory network ($w<0$) balanced with a positive external drive $\mu_{0}>0$. In this case, the system exhibits a unique fixed point as shown in the Appendix, Sec I. If the fluctuations are sufficiently small and the fixed point is stable, the quantity $r_{0}:=F(h_{0},\sigma_{0}^{2})$ then provides an approximation of the mean stationary population activity (or equivalently, the mean stationary firing rate of neurons). At equilibrium, the noiseless version of Eq. (20) gives
$\bar{h}_{0}=\mu_{0}+wr_{0},\quad\sigma_{0}^{2}=\frac{w^{2}(1-p)}{2\tau pN}r_{0},\quad r_{0}=F(\bar{h}_{0},\sigma_{0}^{2})$ (24)
which yields a fixed-point equation for $h_{0}$:
$\frac{\bar{h}_{0}-\mu_{0}}{w}=F\left(\bar{h}_{0},\frac{w(1-p)}{2\tau pN}\left[\bar{h}_{0}-\mu_{0}\right]\right).$ (25)
Note that both sides of the equation correspond to the stationary firing rate $r_{0}$. Furthermore, the fixed points of the 1st-order MF model are obtained by setting $p=1$ in Eq. (25). The solutions for both 1st- and 2nd order MF models can be determined graphically (Fig. 3a): While the left-hand side is identical for the two MF models and is described by a line, the nonlinear function $F$ on the right-hand side of Eq. (25) differs in its second argument $\sigma_{0}^{2}$. For the 1st-order mean-field theory, we have $\sigma_{0}^{2}=0$ and therefore the sigmoidal function is steeper compared to the 2nd-order theory, for which $\sigma_{0}^{2}$ is strictly positive. Because we assume an inhibitory network ($w<0$), the linear function on the left-hand side of Eq. (25) has a negative slope. If we choose $w$ such that the intersection occurs for values $\bar{h}_{0}$ lower than the inflection point $\vartheta=0$ of the hazard function $\phi(h)$ (corresponding to the biologically relevant convex part of $\phi(h)$), then from the graphical solution it is clear that the 2nd-order model exhibits a fixed-point with lower mean input potential $\bar{h}_{0}$ and higher population firing rate $r_{0}$ compared to the 1st-order MF model. This effect becomes stronger for smaller connection probability $p$.
The mean stationary firing rates of the quenched and annealed microscopic models are well captured by the numerical solution of the fixed-point equation (25) (Fig. 3b-d). As a function of the external drive $\mu_{0}$, we observe a strikingly linear dependence between the saturation regimes at low and high values of $\mu_{0}$ (Fig 3b). Furthermore, we find that the firing rate is roughly constant when the connection probability $p$ is varied while $N$ is fixed (Fig 3c). The weak dependence on $p$ is slightly better predicted by the 2nd-order MF theory than by the 1st-order MF theory, which, by definition, has no dependence on $p$. We also observe a power-law behavior of the mean firing rate in both microscopic simulations and fixed-point solutions if the coupling strength is not too strong (Fig. 3d). We note that, in general, the relative error of the 1st-order MF model (and hence the mean-connectivity network) with respect to the stationary mean population rate is rather small compared to the much larger error with respect to the second-order statistics reported in Fig. 1. The second-order statistics will be analyzed later in Sec. IV.5.
IV.1.1 Analytical approximation of fixed points
The linear dependence on $\mu_{0}$, the weak dependence on $p$ and the power-law dependence on $w$ can be explained with our MF theory as follows. Firstly, we notice that for a sufficiently large steepness $\beta\gg 1/\sigma_{0}^{2}$ of the single-neuron transfer function $\phi$, the 2nd-order MF dynamics remains practically unaltered if we take the limit $\beta\to\infty$. In this limit, the single-neuron transfer function tends to $r_{\text{m}}\theta(\bar{h})$, where $\theta$ denotes the Heaviside step function. Indeed, for the values of $\beta$ considered in this paper, taking the limit $\beta\to\infty$ does not noticeably alter the effective nonlinearity $F$ and hence the 2nd-order MF dynamics. The limiting value of the fixed point can be calculated analytically from the graphical representation (Fig. 3) which yields an approximate analytical expression for the firing rate in the 1st-order MF model,
This analytical expression predicts a power-law behaviour as a function of coupling strength in the balanced regime $0<\mu_{0}<-wr_{\text{m}}$. Indeed, the numerical fixed-point solution for the 1st-order MF model exhibits an excellent quantitative agreement with this prediction (Fig. 3d).
Secondly, we note from the geometry of the graphical solution of the fixed points that the intersection point of the 1st-order MF model serves as a rough approximation for the firing rate $r_{0}$ of the 2nd-order MF model as well (Fig. 3a) unless $w$ is extremely negative as discussed below. Indeed, the horizontal stretching of the transfer function in the 2nd-order theory mainly affects the value of the fixed point $\bar{h}_{0}$ rather than the firing rate $r_{0}$. More precisely, a sufficient condition for the validity of our approximation is that $-\mu_{0}/w\gg(1-p)/(\tau C)$ and $-\mu_{0}/w>0.023r_{\text{m}}$ (see Appendix C). We expect that the approximation (26) slightly underestimates the firing rates if the intersection is below the inflection point ($\bar{h}_{0}<0$) and slightly overestimates the firing rates if the intersection is above the inflection point ($\bar{h}_{0}>0$). Our approximation Eq. (26) well explains the piecewise linear behavior of the firing rate as a function of the external input as observed in Fig. 3b. Furthermore, Eq. (26) also explains why the firing rates are roughly independent of the connection probability $p$ (Fig. 3c), and why we see the power-law $-\mu_{0}/w$ as a function of coupling strength $w$ even in non-fully connected networks (Fig. 3d) unless $w$ is extremely negative (this case is treated below). In general we note that, compared to the 1st-order model, the 2nd-order MF model shows a better agreement with the firing rate in microscopic simulations, in particular, it fits the annealed network almost perfectly.
@@ IV.1.1 Analytical approximation of fixed points
IV.1.2 Limit of strongly inhibitory networks
The graphical solution of this equation clearly shows a vanishing firing rate in the 1st-order MF theory ($p\to 1$), whereas the limiting equation has a non-vanishing solution for $p<1$ (2nd-order MF theory, Fig. 3d, inset). The non-vanishing, limiting firing rate in the 2nd-order MF model is in good agreement with simulations of the annealed and quenched microscopic model. In contrast, the 1st-order MF model generally underestimates the firing rates in strongly inhibitory networks: for $w\to-\infty$, the 1st-order MF model has a rate that converges to zero like $-\mu_{0}/w$ as predicted by Eq. (26).
@@ IV.2 Stability of the fixed-point solutions
Figure 4: Second-order MF theory explains stable asynchronous activity and oscillatory instability in the presence of a transmission delay $d>0$. (a) Sample trajectories for quenched (green) and annealed network (red), 2nd-order MF model (blue) and 1st-order MF model (orange) with $d=2$ ms. (b) $(d,w)$-phase diagram with Hopf boundaries (onset of oscillations) for $N=20000$ neurons. The heat map shows the standard deviation of the population rate $r(t)$ in the annealed network to indicate the presence of oscillations. Green line: contour line for the quenched network at which the standard deviation of $r(t)$ equals a threshold of $2$ Hz, which was visually inferred from panel (d). A non-zero threshold was chosen because of finite-size fluctuations. Blue (orange) line: Hopf boundary for 2nd- (1st-) order MF model, calculated from Eq. (31) (c) Standard deviation of the stochastic population rate for different connection strengths $w$ at a delay $d=2$ ms calculated from simulations (circles: quenched network, squares: annealed network, diamonds: 2nd-order MF, triangles: 1st-order MF). (d) Standard deviation of the stochastic population rate for different connection strengths $d$ and fixed coupling strength $w=-1$ mVs. Parameters: $C=100$, $\mu=10$ mV, $\beta=5$ mV-1.
In the presence of a non-zero transmission delay $d>0$, the fixed-point solution of the mean-field model may become unstable, leading to an oscillatory regime, see e.g. [66]. For such an instability to correctly predict oscillatory behavior of our microscopic network model, our mean-field theory must accurately describe the recurrent fluctuations in the network. Comparing simulations of the microscopic network models with simulations of the first- and second-order mean-field models (Fig. 4a), we observe that for the chosen parameters the 1st-order MF model wrongly predicts strong oscillations when the actual population activity of the network is nearly constant corresponding to an asynchronous firing regime. In contrast, the 2nd-order theory correctly reproduces the constant population activity.
@@
To find the onset of oscillations, we look for a Hopf bifurcation, at which $\lambda=i\omega$ with a frequency $\omega\in\mathbb{R}\setminus\{0\}$. Imposing such purely imaginary values of $\lambda$ in Eq. (31) allows us to find the “Hopf boundary” of the oscillatory phase in the parameter space. In the following, we investigate the stability depending on the transmission delay $d$ and the coupling strength $w$ (Fig.  4b). In the $(d,w)$-space, we can see that the 2nd-order MF model significantly differs from the first-order one. For sufficiently strong coupling strength, the 1st-order MF model exhibits a Hopf boundary at small values of $d$, predicting oscillations already at small delays. In contrast, the 2nd-order model predicts oscillation onset for much larger delays. Interestingly, for sparse connectivity (sufficiently small $p$), a stronger coupling strength $w$ requires a larger delay to enable oscillations, whereas the first-order model requires a smaller delay. To assess whether the microscopic models are in an oscillatory state in our simulations, we calculate the standard deviation of a time series of the stochastic population rate $r(t)$ as given in Eq. (6). We find that the oscillatory regime of the annealed network (Fig. 4b, red-colored area) coincides with the area enclosed by the Hopf boundary obtained from the 2nd-order MF theory. A qualitatively similar oscillation boundary is observed in the quenched network model.
For a more detailed comparison of the oscillation amplitude beyond the Hopf bifurcation, we also computed the standard deviation of $r(t)$ in the mean-field models. We find that the 2nd-order MF model matches well the annealed network model (Fig. 4c,d), whereas the 1st-order mean-field model fails to describe the oscillations already at small delays. The quenched microscopic model has an oscillation onset at even larger delays, however, with a qualitatively similar dependence on the delay and coupling strength as the annealed microscopic model.
In conclusion, our stability analysis of the 2nd-order MF model shows that a randomly diluted connectivity leads to a stabilisation of the non-oscillatory state. In particular, the random dilution leads to incoherent fluctuations of the recurrent input, which flattens the effective nonlinearity of the macroscopic dynamics and thus changes the stability properties of the fixed point.

### v2

@@ III.2 Second-order mesoscopic mean-field dynamics
Here, we have used that, for $\sigma=0$, the population transfer function $F(\bar{h},\sigma^{2})$ simplifies to the transfer function of a single Poisson neuron $\phi(\bar{h})=F(\bar{h}(t),0)$ and that $G(\bar{h},0)=0$. Theoretical predictions of MF1, Eq. (21), can always be obtained from the corresponding result of MF2 by taking the limit $p\to 1$ while keeping the network size $N=C/p$ and the coupling strength $w=JpN$ constant.
While the MF1 dynamics, Eq. (21), still captures finite-size effects through the noise term, it does not describe the effect of randomness in the connectivity $\sigma_{w}^{2}>0$ when $p<1$. In fact, MF1 precisely corresponds to the mean-connectivity approximation that underlies the assumption of full connectivity made in previous finite-size theories [42, 43, 47, 36, 44, 45, 46, 33]. This can be seen by applying the diffusion approximation of the Poissonian shot noise for large $N$, Eq. (14), to the exact mesoscopic dynamics, Eq. (7), of the mean-connectivity network, Eq. (5), which recovers MF1. Thus, by comparing simulations of the microscopic models (quenched and annealed networks) with MF1 and MF2, we will thus be able to judge the mean-connectivity approximation used in previous studies [47, 36] and the correction for the effect of random connectivity ($p<1$) given by the 2nd-order MF theory.
In the limit $N\to\infty$, the MF2 theory, Eq. (20), converges to either a one- or a two-dimensional, deterministic dynamics depending on whether the connectivity is dense or sparse. In the dense limit, where $p$ is kept constant and, hence, $C=pN\to\infty$, the variance $\sigma^{2}$ as well as the fluctuations $N^{-1/2}\eta$ and $N^{-1/2}\xi$ vanish, leaving a one-dimensional deterministic firing-rate model. This case is equivalent to the large-$N$ limit of the 1st-order MF theory and can thus neither capture finite-size nor $p<1$ effects. In contrast, in the sparse limit, where $C$ is kept constant and, hence, $p=C/N\to 0$, the noise terms $N^{-1/2}\eta$ and $N^{-1/2}\xi$ vanish but the variance $\sigma^{2}$ does not. The result is a two-dimensional, deterministic mean-field dynamics for the mean and variance of the input potentials,
$\begin{aligned} \tau\frac{\mathrm{d}}{\mathrm{d}t}\bar{h}&=-\bar{h}+\mu(t)+wr(t-d)\\ \tau\frac{\mathrm{d}}{\mathrm{d}t}\sigma^{2}&=-2\sigma^{2}+\frac{w^{2}}{\tau C}r(t-d),\end{aligned}$ (22)
where $r(t)=F(\bar{h}(t),\sigma^{2}(t))$. This sparse connectivity limit is similar to the classical model of Amari [22]. While it captures the effect of random connectivity, it cannot describe finite-size effects and hence second-order statistics caused by finite-size fluctuations. In the following, we will refer to the dynamics in the sparse limit, Eq. (22), in short as MFsparselim.
The effects described by the three population dynamics MF1, MF2 and MFsparselim for different parameter regimes are summarized in Table 3 (Appendix, Sec. J).
@@ III.2.1 Functions $F$ and $G$ for an error-function nonlinearity
being the effective steepness of the sigmoidal nonlinearity at the population level. Because $\beta_{\text{eff}}(\sigma^{2})\leq\beta_{\text{eff}}(0)=\beta$, the effective nonlinearity that governs the population dynamics has a reduced steepness compared to the the single-neuron nonlinearity (Fig. 3a) and this reduction is caused by the strictly positive spread $\sigma>0$ of the membrane potentials in a non-fully connected random network.
@@ IV Analysis of the mesoscopic model
The mesoscopic dynamics MF2, Eq. (20), allows us to calculate and analyze the first- and second-order statistics of mesoscopic quantities such as the stochastic population rate $r(t)$, the population-averaged input potential $\bar{h}(t)$ and the population activity $A_{N}(t,\Delta t)$. We will mainly focus on the second-order statistics, the analysis of which will be based on a linearization of the dynamics about the fixed points of the noiseless systems. Therefore, we will first calculate the fixed points and their stability. Although not the focus of the paper, this analysis will also yield approximations for first-order statistics including the stationary mean population activity and the linear response of the mean population activity to time-dependent external stimuli.
@@ IV.1 Stationary mean population activity
We first calculate the fixed-points $\bar{h}(t)=\bar{h}_{0}$, $\sigma^{2}(t)=\sigma_{0}^{2}$ and $\xi(t)=0$ of the deterministic system when the noise terms in Eq. (20) are set to zero and the external drive is constant, $\mu(t)=\mu_{0}$. In the following, we assume an inhibitory network ($w<0$) balanced with a positive external drive $\mu_{0}>0$. In this case, the system exhibits a unique fixed point as shown in the Appendix, Sec. I. If the fluctuations are sufficiently small and the fixed point is stable, the quantity $r_{0}:=F(h_{0},\sigma_{0}^{2})$ then provides an approximation of the mean stationary population activity (or equivalently, the mean stationary firing rate of neurons). At equilibrium, the input potential $h_{0}$ is given by the fixed-point equation
$\frac{\bar{h}_{0}-\mu_{0}}{w}=F\left(\bar{h}_{0},\frac{w(1-p)}{2\tau pN}\left(\bar{h}_{0}-\mu_{0}\right)\right),$ (25)
(see appendix, Sec. B). Furthermore, the fixed points of the MF1 are obtained by setting $p=1$ in Eq. (25) corresponding to setting $\sigma_{0}^{2}=0$ in the sigmoidal function $F(h_{0},\sigma_{0}^{2})$. The solutions for both MF1 and MF2 can be determined graphically (Fig. 3a): For MF1, $\sigma_{0}^{2}=0$ and therefore the sigmoidal function $F(h_{0},0)$ is steeper compared to MF2, for which $\sigma_{0}^{2}$ is strictly positive.
While the fixed-points can be obtained numerically, for the analytical theory we use the approximations
$r_{0}$ $\approx-\frac{\mu_{0}}{w},$ (26)
$\bar{h}_{0}$ $\approx\Phi^{-1}\left(-\frac{\mu_{0}}{wr_{\text{m}}}\right)\sqrt{\beta^{-2}+\sigma_{0}^{2}}$ (27)
@@
for stimuli $0<\mu_{0}<-wr_{\text{m}}$, where the firing rates are low ($0<r<r_{\text{m}}/2$), see appendix, Sec. B. The approximation holds for large $\beta\gg 1/\sigma_{0}^{2}$ and for $\mu_{0}/w\gg(1-p)/(\tau C)$ and $\mu_{0}/w>0.023r_{\text{m}}$. The fixed-point solutions yield theoretical predictions for the stationary first-order statistics such as the stationary mean firing rate $\left\langle r\right\rangle$, mean membrane potential $\left\langle\bar{h}\right\rangle$ and mean membrane-potential spread $\left\langle\sigma^{2}\right\rangle$. However, we postpone the discussion of these predictions to the Appendix B, because our main focus is the second-order statistics.
@@ IV.2 Stability of the fixed-point solutions
Figure 3: Analysis of the deterministic mean-field dynamics. (a) Graphical solution of the fixed point equation (25), here illustrated for $\mu_{0}=5$ mV and w=0.7w=-0.7 mVs. Solid lines: function $F(\bar{h},\sigma^{2}(\bar{h}))$ for MF1 (orange, $\sigma^{2}(\bar{h})\equiv 0$) and MF2 (blue, $\sigma^{2}(\bar{h})$ as in Eq. (25)). Black dashed line: left-hand side of Eq. (25). Cyan dash-dotted line: $F$ with the second argument fixed at the fixed-point value $\sigma_{0}^{2}$. (b) $(d,w)$-phase diagram with Hopf bifurcation lines (onset of oscillations). Bifurcation lines of MF2 and MFsparselim explain stable asynchronous activity and oscillatory instability. The heat map shows the standard deviation of the population rate $r(t)$ in the quenched network ($p=0.1$) to indicate the presence of oscillations. Hopf bifurcations, calculated from Eq. (30), are displayed for MF1 (orange line), MF2 (blue line) and MFsparselim (brown line). (c) Sample trajectories of population rate $r(t)$ for quenched network (green), MF1 (orange), MF2 (blue) and MFsparselim (brown) with $d=1$ ms. (d)-(f) Linear response of the time-dependent mean population rate (“rate susceptibility”) to a weak modulation of the external current $\mu(t)$ for quenched network (green circles), MF1 (orange), MF2 (blue) and MFsparselim (dashed brown). The MF theories are given by Eq. (34), where $p$ is set to $1$ for MF1, and the sparse limit $N\to\infty$, $pN=C=\text{const}$ is taken for MFsparselim. For each panel, top and bottom plots show the amplitude $|\tilde{\chi}_{r}|$ and the phase $\phi_{r}$ of the susceptibility, respectively. Dashed horizontal line: $\phi_{r}=-\pi/2$. (d) $p=0.1$ ($C=100$), $d=0$; (e) $p=0.4$ ($C=400$), d=1d=1\,ms; (f) $p=0.95$ ($C=950$), $d=0$. Other parameters unless otherwise indicated: $N=1000$, w=1w=-1 mVs, $\beta=5$ mV-1, $\mu_{0}=10$ mV, $\sigma_{\text{ext}}=0$.
In the presence of a non-zero transmission delay $d>0$, the fixed-point solution of the mean-field model may become unstable, leading to an oscillatory regime, see e.g. [73]. For such an instability to correctly predict oscillatory behavior of our microscopic network model, our mean-field theory must accurately describe the recurrent fluctuations in the network.
@@
To find the onset of oscillations, we look for a Hopf bifurcation, at which $\lambda=i\omega$ with a frequency $\omega\in\mathbb{R}\setminus\{0\}$. Imposing such purely imaginary values of $\lambda$ in Eq. (30) allows us to find the “Hopf boundary” of the oscillatory phase in the parameter space. In the following, we investigate the stability depending on the transmission delay $d$ and the coupling strength $w$ (Fig.  3b). In the $(d,w)$-space, we can see that the bifurcation lines of the noiseless systems significantly differ between 2nd-order and first-order MF dynamics. For sufficiently strong coupling strength, the 1st-order MF dynamics exhibits a Hopf boundary at small values of $d$, predicting oscillations already at small delays. In contrast, the 2nd-order theory predicts an oscillation onset for much larger delays. Interestingly, MF2 predicts that a stronger coupling strength $w$ requires a larger delay to enable oscillations, whereas the first-order model requires a smaller delay. We note that the bifurcation line obtained for MFsparselim is close to MF2. The bifurcation lines qualitatively explain the presence or absence of large noisy oscillations in the mean-connectivity and random network models (Fig.  3b,c). Furthermore, the simulations of the MF dynamics indicate that the Hopf bifurcation is supercritical: varying the delay, the variance of the population rate $r$ changes continuously at the bifurcation and no bistability between fixed point and limit cycle is observed.
In conclusion, our stability analysis of MF2 shows that connectivity disorder leads to a stabilization of the non-oscillatory stationary state. In particular, the disordered connectivity leads to incoherent fluctuations of the recurrent input, which flattens the effective nonlinearity of the macroscopic dynamics and thus changes the stability properties of the fixed point.